Solve by the substitution method
2x + 7 =12
x = 8 - 4y

Answer by nerdybill(7384)

2x + 7 =12 <--equation 1
x = 8 - 4y <--equation 2
.
Notice equation 2 states that'x' is equal to (8 - 4y) -- substitute this into equation 1:
2x + 7 =12
2(8 - 4y) + 7 =12
16 - 8y + 7 =12
23 - 8y =12
-8y = -11
y = 11/8 (solution for y)
.
substitute the above into equation 2 and solve for x:
x = 8 - 4y
x = 8 - 4(11/8)
x = 8 - 11/2
x = 16/2 - 11/2
x = 5/2 (solution for x)